OpenStudy (anonymous):

1/5 log(base 2)5 = ?

OpenStudy (anonymous):

if its like this (1/5) log(base 2)5 = ? by changing the base (1/5) (log5/log 2) =0.464385619 if its like this 1/[5 log(base 2)5] = =1/[5(log5)/(log2)] =0.0861353